OpenStudy (anonymous):

find limit as t approaches 0 of sin 3t/2t

OpenStudy (anonymous):

hint \[\frac{\sin(3t)}{2t}=\frac{\sin(3t)}{3t}\times \frac{3}{2}\]

OpenStudy (anonymous):

thanks!! :)